4. The lateral edge of a square prism is 20 cm long and it is inclined at an angle of 60 degrees with the base. If the volume of the prism is 1,200 cm3, determine the base edge.
1
Expert's answer
2020-12-10T13:30:12-0500

Consider a regular square pyramid

Let "a=" the base edge, "h=" the altitude, and "l=" the lateral edge.

The volume "V" of the of a regular square prism is


"V=a^2h"

Then

"a=\sqrt{\dfrac{V}{h}}"


From the right triangle


"h=l\sin60\degree"

"a=\sqrt{\dfrac{1200}{20(\dfrac{\sqrt{3}}{2})}}=2\sqrt{10\sqrt{3}}(cm)"

"a=2\sqrt{10\sqrt{3}}\ cm"